# Guided Multi-Fidelity Bayesian Optimization for Data-driven Controller Tuning with Digital Twins

### v1

@@ V-A Initial Dataset Ablation Study
$N_{0_{s=1}}$} $N_{0_{s^{\prime}}}$} EI-BO MFBO GMFBO
@@
V-B Comparison with Baseline Methods
We compare GMFBO against baseline BO with Expected Improvement (EI) and three MFBO variants: baseline MFBO with trace-aware Knowledge Gradient (taKG) [35], baseline MFBO with our caEI, and a modified MFBO with caEI using the fidelity-aware kernel in (13) without IS3. The taKG acquisition weighs expected information gain against sampling cost, which relies strongly on the surrogate GP. GMFBO explicitly corrects DT bias and guides optimization toward the objective. In contrast, taKG resembles an active learning strategy that improves surrogate accuracy in informative regions without adapting to biased or shifting fidelities.
Figure 5: Unbiased sampling cost over BO iterations, averaged over $50$ Monte Carlo experiments. Shaded regions indicate $95\%$ confidence intervals.
To compare sampling costs across IS selections, we define the unbiased sampling cost of an experiment with $n$ data points as $\sum_{i=1}^{i=n}\mathcal{H}(s_{i},\hat{e}_{\text{IS2}})-b(\hat{e}_{\text{IS2}})=\sum_{i=1}^{i=n}s_{i}$, which quantifies the number of samples acquired from each information source. Figure 5 shows the mean and confidence intervals (CI) of these costs for different methods. GMFBO has a higher initial sampling cost since $N_{0_{s=1}}\cdot n_{\text{c}}=2\cdot 4=12$ IS3 samples are required. Since IS3 is inexpensive and guides the optimizer to reduce IS1 queries, GMFBO attains a lower cumulative unbiased cost than BO and other MFBO methods, whose higher variance results from oversampling inaccurate DT data, inflating the sampling cost.
@@ V-B Comparison with Baseline Methods
Fig. 6 shows that baseline MFBO with taKG (olive) fails to exploit IS2 and exhibits higher variance than MFBO with caEI (purple) due to poor acquisition. Large IS2 errors cause taKG to oversample inaccurate IS2 data, polluting the surrogate without adjusting fidelity influence. Since taKG estimates the optimum from the posterior GP, its failure arises from IS2’s relative absolute error (Fig. 4(b)), which exceeds $20\%$ on average. Using caEI improves the baseline MFBO to a level comparable to that of BO using EI, although without a clear reduction in CI. The modified MFBO (black) further benefits from the customized kernel, which shrinks variance by adjusting the influence of DT data in the surrogate and preventing optimizer pollution from low-accuracy DR data, leading to improved convergence. Our GMFBO, which combines caEI, a customized kernel, and IS3 correction, achieves higher data efficiency by requiring fewer IS1 measurements with reduced and faster-decreasing variability across iterations. BO and modified MFBO with caEI need $22$ and $11$ IS1 experiments on average (including initialization) to reach a given threshold, whereas GMFBO requires only $6$, i.e., $72\%$ and $45\%$ fewer, respectively. In the best run, GMFBO converges in $2$ experiments, compared to $8$ and $6$ for BO and modified MFBO, yielding $75\%$ and $66\%$ improvements. Even in the worst case, GMFBO converges within $8$ experiments, while BO may not converge after $22$, and modified MFBO requires $11$, i.e., improvements of at least $64\%$ and $27\%$. These results show that GMFBO outperforms BO and both baseline and modified MFBO methods.
V-C Adaptation to System Variation
@@
As shown in Fig. 7, GMFBO adapts to the IS1 change and finds the new optimum in $10$ IS1 experiments, compared to over $18$ for BO with EI ($44\%$ improvement). The change increases IS2 error, which lowers cross-source correlation and raises IS2 sampling cost, while online IS3 corrections improve the surrogate. This allows GMFBO to continue exploiting DT data despite shifting system dynamics.
VI Hardware Experimental Results
We validate data efficiency and generalization on the Maxon HEJ 90 drive system (Fig. 8), with IS1 as the real hardware and IS2 as the DT (corresponding to IS1 in the previous section V). The integrated EPOS4 controller employs a hierarchical structure comprising current, position, velocity, and torque loops, all of which are accessed via EtherCAT. A custom MasterMACS–ApossIDE integration enables real-time acquisition of encoder positions and control inputs at $1$ kHz [36].
@@ VI Hardware Experimental Results
Figure 9: Sampled data points of GMFBO and baseline BO with EI in a single experiment. Iterations are shown until the minimum objective approaches the global optimum (star). Contours depict the true objective.
Fig. 9 shows sampled data points per method in a single experiment until the optimum is found. Baseline BO relies heavily on costly IS1 measurements, whereas GMFBO shifts part of the exploration and exploitation to cheaper IS2 and IS3, reducing IS1 usage.
Figure 10: Unbiased sampling cost over BO iterations from 2020 Monte Carlo experiments. Shaded regions denote $95\%$ confidence intervals.
We conduct a Monte Carlo analysis with $N_{\text{exper}}=20$ experiments. Fig. 10 shows that our GMFBO method incurs lower total sampling cost than both BO and MFBO over $N=20$ BO iterations. We also observe variability in the unbiased sampling cost for GMFBO and MFBO, stemming from the dynamic selection between two information sources guided by our acquisition function. Additionally, the confidence interval for GMFBO’s sampling cost is wider than that of MFBO, reflecting greater cost variability due to reliance of IS3 to the real system measurements.
@@
Fig. 11 shows that GMFBO requires fewer real-system experiments on average, with a smaller and faster-shrinking confidence interval than other methods. BO and MFBO need $9$ and $7$ experiments, respectively, while GMFBO achieves the same performance in only $5$, yielding $44\%$ and $28\%$ improvements. Among the experiments, the best GMFBO converges in $2$ experiments (vs. $6$ for BO and $4$ for MFBO, $67\%$ and $50\%$ improvements), and the worst in $6$ (vs. $13$ and $8$, $54\%$ and $25\%$ improvements).
@@ VII Conclusion
These results demonstrate the potential of GMFBO to accelerate learning in systems with costly experiments and imperfect models. Future work will extend the framework to settings where fidelity varies with the controller gains (input-dependent) or the system’s operating conditions (state-dependent, e.g., velocity or load). We also plan to investigate closed-loop, data-efficient digital twin identification and robustness under non-convex and discontinuous objectives.

### v2

@@ V-A Comparison with Baseline Methods
$N_{0_{s=1}}$} $N_{0_{s=0}}$} EI-BO MFBO (taKG) GMFBO
@@
To compare sampling costs across IS selections, we assign a cost of $1$ to each IS1 query and $0.1$ to each IS2 or IS3 query, reflecting the fact that simulations are an order of magnitude cheaper than real-system evaluations while still contributing informative samples. Figure 5 shows the mean and confidence intervals (CI) of these costs for different methods. GMFBO has a higher initial sampling cost since $N_{0_{s=1}}\cdot n_{\text{c}}=2\cdot 4=8$ IS3 samples are required. Since IS3 is inexpensive and guides the optimizer to reduce IS1 queries, GMFBO attains a lower cumulative sampling cost than BO and other MFBO methods, whose higher variance results from oversampling inaccurate DT data, inflating the sampling cost.
Figure 5: Sampling cost over BO iterations, averaged over $50$ Monte Carlo experiments. Shaded regions indicate $95\%$ confidence intervals.
Figure 6 shows that baseline MFBO with taKG (olive) fails to exploit IS2 and exhibits higher variance than MFBO with caEI (purple) due to poor acquisition. Large IS2 errors cause taKG to oversample inaccurate IS2 data, thereby distorting the surrogate model due to the lack of fidelity-aware weighting. Since taKG estimates the optimum from the posterior GP, its failure arises from IS2’s relative absolute error $\frac{\lvert\mathsf{g}_{\text{c}}(k,1)-\mathsf{g}(k,0)\rvert}{\lvert\mathsf{g}_{\text{c}}(k,1)\rvert}$ visualized in Figure 4, which exceeds $20\%$ on average. Using caEI improves the baseline MFBO to a level comparable to that of BO using EI, although without an apparent reduction in CI. The modified MFBO (black) further benefits from the customized kernel, which shrinks variance by adjusting the influence of DT data in the surrogate and preventing optimizer pollution from low-accuracy DR data, leading to improved convergence. Our GMFBO, which combines caEI, a customized kernel, and IS3 correction, achieves higher data efficiency by requiring fewer IS1 measurements with reduced and faster-decreasing variability across iterations. BO and modified MFBO with caEI need $22$ and $11$ IS1 experiments on average (including initialization) to reach a given threshold, whereas GMFBO requires only $6$, i.e., $72\%$ and $45\%$ fewer, respectively. In the best run, GMFBO converges in $2$ experiments, compared to $8$ and $6$ for BO and modified MFBO, yielding $75\%$ and $66\%$ improvements. Even in the worst case, GMFBO converges within $8$ experiments, while BO may not converge after $22$, and modified MFBO requires $11$, i.e., improvements of at least $64\%$ and $27\%$. These results show that GMFBO outperforms BO and both baseline and modified MFBO methods.
@@
V-B Adaptation to System Variation
@@
As shown in Figure 7, GMFBO adapts to the IS1 change and finds the new optimum in $10$ IS1 experiments, compared to over $18$ for BO with EI ($44\%$ improvement). The change increases IS2 error, which lowers cross-source correlation and raises IS2 sampling cost, while online IS3 corrections improve the surrogate. This allows GMFBO to continue exploiting DT data despite shifting system dynamics.
VI Hardware Validation
We validate data efficiency and generalization on the Maxon HEJ 90 drive system (Figure 8), with IS1 as the real hardware and IS2 as the DT (corresponding to IS1 in Section V). The integrated EPOS4 controller employs a hierarchical structure comprising current, position, velocity, and torque loops, all of which are accessed via EtherCAT. A custom MasterMACS–ApossIDE integration enables real-time acquisition of encoder positions and control inputs at $1$ kHz [36].
@@ VI Hardware Validation
Figure 9: Sampled data points of GMFBO and baseline BO with EI in a single experiment. Iterations are shown until the minimum objective approaches the global optimum (star). Contours depict the true objective for the real hardware.
Figure 9 shows sampled data points per method in a single experiment until the optimum is found. Baseline BO relies heavily on costly IS1 measurements, whereas GMFBO shifts part of the exploration and exploitation to cheaper IS2 and IS3, reducing IS1 usage.
Figure 10: Sampling cost over BO iterations from 2020 Monte Carlo experiments. Shaded regions denote $95\%$ confidence intervals.
We conduct a Monte Carlo analysis with $N_{\text{exper}}=20$ experiments. Figure 10 shows that our GMFBO method incurs lower total sampling cost than both BO and MFBO over $N=20$ BO iterations. We also observe variability in the sampling cost for GMFBO and MFBO, stemming from the dynamic selection between two information sources guided by our acquisition function. Additionally, the confidence interval for GMFBO’s sampling cost is wider than that of MFBO, reflecting greater cost variability due to IS3’s reliance on real system measurements.
@@
Figure 11 shows that GMFBO requires fewer real-system experiments on average, with a smaller and faster-shrinking confidence interval than other methods. BO and MFBO need $9$ and $7$ experiments, respectively, while GMFBO achieves the same performance in only $5$, yielding $44\%$ and $28\%$ improvements. Among the experiments, the best GMFBO converges in $2$ experiments (vs. $6$ for BO and $4$ for MFBO, $67\%$ and $50\%$ improvements), and the worst in $6$ (vs. $13$ and $8$, $54\%$ and $25\%$ improvements).
@@ VII Conclusion
These results demonstrate the potential of GMFBO to accelerate learning in systems with costly experiments and imperfect models. Future work will extend the framework to settings where fidelity varies with the controller gains (input-dependent) or the system’s operating conditions (state-dependent, e.g., velocity or load). Other promising directions are closed-loop, data-efficient digital twin identification and robustness under non-convex and discontinuous objectives.